OpenStudy (anonymous):

Find a formula for f(x) f(2) = (1/2)^0 +(1/2)^1 + (1/2)^2 = 1+ 1/2+1/4=7/4

OpenStudy (anonymous):

What is the original function?

OpenStudy (anonymous):

I can't understand the question

OpenStudy (anonymous):

let f(x) be a function that adds the first x powers of 1/2 and then find the function with this given info.

OpenStudy (anonymous):

Let f(x) be a function that adds the first x powers of 1/2: f(x) = (1/2)^0 = 1 f(1) = (1/2)^0 + (1/2)^1 = 1+1/2 =3/2 f(2) = (1/2)^0 +(1/2)^1 + (1/2)^2 = 1+ 1/2+1/4=7/4

OpenStudy (kirbykirby):

Is this supposed to go on to infinity? i.e \[\sum_{x=0}^{\infty}\left( \frac{1}{2}\right)^x \]

OpenStudy (anonymous):

how do i express that in terms of f(x) ?

OpenStudy (kirbykirby):

maybe you could use a recursive definition \( f(x_0)=1/2\\ f(x_n)=f(x_{n-1})+1/2\)